Prove by contradiction that for any integer n if n2 is odd then n is odd.


1
Expert's answer
2021-10-11T12:20:10-0400

Consider the next statement: for any integer if "n^{2}" is odd then n is non-odd (if n is non-odd integer then n is even)

So, if "n^{2}" is odd then n is even

Let "n^{2}" = 2k+1, k "\in" Z

n = 2m, m "\isin" N ∪ {0}

"n^{2} = 2k+1 \to n=\sqrt{2k+1}"

"\sqrt{2k+1} = 2m \to 2k+1=4m^{2}"

Since 2k is even integer, then 2k+1 is odd. "4m^{2}" is also even. We came to the conclusion that an odd number is equal to the even integer number, which is false. So, assumption that ' for any integer if "n^{2}" is odd then n is non-odd ' is false, which means that for any integer if "n^{2}" is odd then n is odd. The statement has been proved.